OpenStudy (anonymous):

SQUAREROOT OF 225 + SQUARE ROOT OF 361 CAN ANYONE EXPLAIN? I KNOW THE ANSWER ANY SHORTCUT METHODS??

OpenStudy (katrinakaif):

I just memorize the roots till up to 20..=/ That'd be helpful.. o_o

OpenStudy (anonymous):

Yep, no mathematical shortcuts unfortunately, just have to work out the square roots first.

OpenStudy (anonymous):

OK THANKS :D

OpenStudy (anonymous):

I M NOT THAT GOOD IN MEMORIZING :(

OpenStudy (anonymous):

1^2=1 2^2=4 3^2=9 4^2=16 5^2=25 6^2=36 7^2=49 8^2=64 9^2=81 10^2=100 11^2=121 12^2=144 13^2=169 14^2=196 15^2=225 16^2=256 17^2=289 18^2=324 19^2=361 20^2=400 : Memorize this ... It will help u a lot..